Given that y= acoskx + bsinkx, show that d²y/dx² + k²y =0

1
Expert's answer
2022-03-28T04:18:43-0400

"y(x)=acos(kx)+bsin(kx)"

"y'(x)=-a\times ksinkx+b\times kcos(kx)"

"y"(x)=-a\times k^2\times cos(kx)-b\times k^2sin(kx)"


"=-k^2(acos(kx)+bsin(kx))"

From this its clear that,

"y"(x)=-k^2\times y(x)"

y"+k2y=0


Hence,

"\frac{d^2y}{dx^2}+k^2y=0"